600 lb 600 600 lb 300 lb 6 ft 300 lb 6 ft

Using the method of joints, determine the force in each member.
Answer #1

Sol nom the Groma Goolb Goolb tan e= 6 AYAe el -Gos (6) - 600 (ic) Go(24) 300 (32)+) F 1200 1lb Au t Hy-3(600) _ 2(30): o AyFag cas 36.87十FAC-O - 15D0 x cos 36.87FAc FAC = 1200 lb (t) S H 个FBC FAC 600lb FBD 36-87 36.87 Fee A 8FBD-FBE =-500(2) 2FBD- 2.000 oh e Fee = -500 lb Fee FEF = 500 lb (c) Joint E FEF ED -Fee gin 36.87 36.87 Eq -FEF ging 6.872